OpenStudy (anonymous):

The expression (2x)-3 can be rewritten as

OpenStudy (ybarrap):

2x-3, you sure that's it?

OpenStudy (anonymous):

(2X) To the -3 power

OpenStudy (raden):

a^-n = 1/a^n

OpenStudy (raden):

therefore (2x)^-3 = 1/(2x)^3 = 1/(2^3 * x^3) = 1/(8x^3)

OpenStudy (anonymous):

thank u

OpenStudy (raden):

you're welcome